The following payoff table provides profits based on various possible decision alternatives and various levels of demand at Amber Gardner's software firm:

Demand Level
0.3 0.7
Low High
Alternative A $10,000 $30,000
B $5,000 $40,000
C ($2,000) $50,000
*Profits in $ thousands

a. Plot the expected-value lines on a graph. (Answered below)

Alternative Demand Level
0 1
A $ 10,000.00 $ 30,000.00
B $   5,000.00 $ 40,000.00
C $ (2,000.00) $ 50,000.00

560,000.00 550,000.00 $40,000.00 530,000.00 520,000.00 $10,000.00 0 S(10,000.00)

b. Is there any alternative that would never be appropriate in terms of maximizing expected profit? Explain on the basis of your graph. (see graph in part a)

c. For what range of P(High Demand) would alternative A be the best choice if the goal is to maximize expected profit?

d. For what range of P(High Demand) would alternative B be the best choice if the goal is to maximize expected profit?

e. Compute the expected values for each alternative if the probability of low demand level is 0.30. Which of the the options is best under this probability?

f. Using the probability of low demand as 0.3 (therefore probability of high demand = 0.7), compute the EVPI. Explain the significance of this number (i.e. what does it mean?).
Answer #1

a)

1) Expected profit Probability tligg 2 Low 0 4 0.05 5 0.1 -10000 A4+B4 30000 5000*A4+40000 B4 -2000 A4+50000 B4 =1-A4 1-A5 =

G23 0 Probability 2 LowHigh A Expected profit 30000 0 2900038250 47400 0.9 0 2600033000 0.75 0.7 0 0.6 0.55 0.4 0.45 13 05 50

b)

There is no such alternative. Every alternative has the highest payoff at some probability range

c)

Alternative A is best for H=0 and above

Let the upper limit be p

When P(high) =p, A and B would be same and after that B would give better payoff

To find p, Expected payoff of A = expected payoff of B

10000*(1-p)+30000*p = 5000*(1-p)+40000*p

p = 0.333

d)

The lower limit is found is part c.

At the upper limit, ::

At this probability payoff of B and C would be same

Let probability of high demand be p

5000*(1-p)+40000*p = -2000*(1-p)+50000*p

p = 0.412

Therefore, B is best for 0.333<= P(high) <= 0.412

e)

1) Expected profit 2 Low High A 0.3 0.7 24000 29500 34400

Can be seen from the table in the first part

f)

Expected value with out perfect information = max(24000,29500,34400) = 34400

When there is perfect information, under each demand condition, best possible alternative is chosen.

Expected value with perfect information = 0.3*10000+0.7* 50000 = 38000

EVPI = 38000-34400= 3600